OpenStudy (anonymous):

find the inverse of 6^(x-5)

OpenStudy (anonymous):

y = 6^(x-5) , switch x and y x=6^(y-5) , now solve for y

OpenStudy (anonymous):

your inverse is x = 6^(y-5)

OpenStudy (anonymous):

my teacher answer is f-1(x)=5+log6x i dont know how she got it

OpenStudy (anonymous):

solve for y satan

OpenStudy (anonymous):

satan?

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

i am a satan worshipper

OpenStudy (anonymous):

interesting but im not sure how to solve for y can you walk me through

OpenStudy (anonymous):

i just prayed for satan's help

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

take that kind of talk somewhere else, you may be playing but we are here to solve math problems

OpenStudy (anonymous):

^^what lagrange said but i dont judge

OpenStudy (anonymous):

hey, dont persecute my religion

OpenStudy (anonymous):

ok x = 6^(y-5)

OpenStudy (anonymous):

take log base 6 of both sides

OpenStudy (anonymous):

can you demonstrate how you were earlier with the numbers

OpenStudy (anonymous):

satan ish baladdama log_6 (x) = log_6 ( 6^(y-5)

OpenStudy (anonymous):

first i switched x and y

OpenStudy (anonymous):

yea so x= 6^y-5 right?

OpenStudy (anonymous):

yes, now take log base 6 of both sides

OpenStudy (anonymous):

that will cause the right side to cancel

OpenStudy (anonymous):

log6x= log6y-5?

OpenStudy (anonymous):

is that what you mean by take log 6 of both sides?

OpenStudy (anonymous):

you have to be more careful

OpenStudy (anonymous):

log_6 (x) = log_6 [ 6^(y-5)]

OpenStudy (anonymous):

but you have the right idea ;)

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

the right side should cancel

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so after the right side cancels we are left with log_6 (x) = log_6 [ 6^(y-5)] log_6 (x) = (y-5)

OpenStudy (anonymous):

ok what do you do next

OpenStudy (anonymous):

now add 5 to both sides

OpenStudy (anonymous):

log_6 (x) = log_6 [ 6^(y-5)] log_6 (x) = (y-5) log_6 (x) = log_6 [ 6^(y-5)] log_6 (x)+5 = y

OpenStudy (anonymous):

log_6 (x) = log_6 [ 6^(y-5)] log_6 (x) = (y-5) log_6 (x)+5 = y